8. ABC has taken up a turnkey contract which is valued at INR 600 Million, to be executed over a period of 6 months ABC erpects to spend 400 Million as total However, in the very first month of the project, steel prices went up consideraby were to be purchased in the first month for 70 Million were actually purchased for i too with a delay of 1 month as the procurement managers manage the overshooting costs. Other costs in the first two months were as per costs for executing the project and items which 00 Million;, that tried to do opportunistie buying to plan As per plan, items worth 150 Million are still to be purchased (which will be done in the last mont the project, and the PM estimates that these will end up costing 20% higher than planned The delay in the project is likely to cause one additional month of working, which will lead to additional manpower costs of 20 Million. The PM estimates that other than these two factors, there will be no other impact of costs in the remaining months The PM is happy to note that there is no financial penalty for the project delay and the client w make the full payment of 600 Million after the project closure. At the end of 2 month the following is the status of the project - Revenues Billed to client: 300 Million INR Payment Made by the client 200 Million INR - Actual Costs of activities completed and procurement done: 120 Million - Costs of activities completed, but vendor invoices still not received: 20 Milion Analyze the Financial Performance of the project, and report the following (8 marks) a Baseline Planned Profitability (as a %) b. Current profitability (as a %)
Answer #1

a)

Total revenue expected from the project at the end of 6 months = $ 600 Mn

Total costs expected to be incurred = $ 400Mn

Therefor, the Expected profits are = $ 600Mn - $ 400 Mn = $ 200 Mn

Baseline profitability = 200/600 = 33.33%

b)

Since, as per US GAAP firms need to do accrual accounting, accrued revenues and costs will be used for calculating profitability.

Accrued revenue is what the firm has billed to the client and accrued costs is the cost of all activities completed.

So, revenue = $300 Mn

Costs = $ 180 Mn + $ 20 Mn = $ 200 Mn

Therefore, the current profitability = (300-200)/300 = 100/300 = 33.33%